# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

### v2

@@ Acknowledgements
We extend our sincere gratitude to Viktoria Schram for providing assistance to reproduce baselines.
Appendix A Experimental Details
@@
   FLoRes-200 Costa-jussà et al. (2022b) afr, amh, arz, bak, bel, ceb, cym, dik, ewe, fao, Multi-domain
guj, hau, hne, hye, ibo, ind, jav, kan, kat, kaz,
khm, kin, kir, lmo, ltz, mar, mri, mya, oci, pan,
plt, sin, sna, snd, som, ssw, tam, tat, tgl, tuk,
wol, xho, yor, zul
@@
      Neulab TED Talk Qi et al. (2018) bel, hye, kat, kaz, mar, mya, tam TED
      News Commentary kaz Economics
      OPUS100 Zhang et al. (2020) afr, amh, bel, cym, guj, hau, hye, ibo, ind, kan, Multi-domain
kat, kaz, khm, kin, kir, mar, mya, oci, pan, sin,
tam, tat, tuk, xho, yor, zul
      OPUS Bible afr, amh, ceb, dik, ewe, guj, hye, ind, kan, mar, Religion
      Christodouloupoulos and Steedman (2015) mri, mya, plt, sin, sna, som, ssw, tam, tgl, wol,
xho, zul
      OPUS OpenSubtitles afr, hye, kat, sin, tgl Movies
@@
ind, jav, kan, kat, kaz, khm, kin, kir, lmo, ltz,
mar, mri, mya, pan, sna, tat, tgl, tuk, yor, zul
      OPUS Tanzil Tiedemann (2012) amh, hau, ind, snd, som, tat Religion
      OPUS Gnome Tiedemann (2012) fao, mri, som Technology
      OPUS GlobalVoices Tiedemann (2012) amh, mya News
@@
      WikiMatrix Schwenk et al. (2021) arz, bak, bel, ceb, fao, ind, jav, kat, kaz, lmo, Wiki
ltz, mar, oci, sin, tam, tat, tgl
@@
Table 7: List of datasets under study covering 50 different languages. We only opt for 44 out of 500 languages available in the MT560 dataset, 44 out of 200 languages available in the FLoRes-200 dataset, and 8 out of 12 languages available in the NusaTranslation dataset.
A.1 Languages Under Study
We list all the languages used in the training from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets in Table 8 and Table 9, respectively. The language code follows ISO639-3 coding. All languages are also complemented by their rarity taxonomy based on Joshi et al. (2021) into two vitality classes: 0-2→low resource language (LRL), and 3→mid resource language (MRL). We also provide information about whether the language was part of the pre-trained M2M100 model dataset to highlight the model knowledge coverage.
A.2 Models
Here are the details on the proxy LMs we use in the experiments as follows:
@@ A.2 Models
  • NLLB (No FT) Costa-jussà et al. (2022b): [^4] NLLB (1.3B) is taken from https://github.com/facebookresearch/fairseq/tree/nllb. a pre-trained estimated model of NLLB-200 Distilled (1.3B) without any fine-tuning. We run the model in a zero-shot fashion.
Language Language Code Family Joshi Class Vitality Seen by M2M100
@@
Table 8: List of languages from the English-centric dataset, including their rarity category mapping and an indication of whether they are involved in the pre-training process for M2M100. Note that all languages are involved in the pre-training process for NLLB.
A.3 Hyper-parameters
@@
Each fine-tuning and evaluation for LMs is done with an NVIDIA Tesla V100 32GB GPU. The hyper-parameters used during fine-tuning from the English-centric and Many-to-Many Languages datasets are listed in Table 10, 11, 12, and 13 for SMaLL100, M2M100, NLLB, and Transformer models, respectively.
@@
Each regressor is trained on an AMD Ryzen Threadripper 2990WX with 128 GB of RAM and 16 threads. Regressors’ hyper-parameters used are provided in Table 14, 15, 16, and 17 for XGBoost, Poly2/Poly3, LGBM, and MF, respectively. These hyper-parameters were obtained based on the best cross-validation RMSE score using 10 folds.
@@
Table 9: List of languages from the Many-to-Many Languages dataset along with their rarity category mapping and an indication of whether they are included in the pre-training process for each respective model.
@@
A.4 Regressor Dataset Sizes
@@
Appendix B More Detailed Results
We provide detailed visualizations of the results of XGBoost with ProxyLM Ensemble based on multiple language groupings in Figures 5-10 for English-centric result, and Figures 11-14 for Many-to-Many Languages result. Each language groupings plot comprises multiple subplots, including (a) vitality class, (b) Joshi class, (c) language family, and (d) individual languages. The mapping of vitality, Joshi class, and language family follows the classifications in Table 8 and 9.
Appendix C Feature Importances

### v3

@@ Acknowledgements
We extend our sincere gratitude to Viktoria Schram for providing assistance in reproducing Matrix Factorization as our baseline.
Appendix A Dataset Features Calculation
In this section, we will describe in detail how the dataset features were computed.
  • Train size: The number of dataset entries used for training specified NLP task.
  • Vocab size: The number of unique tokens in the dataset after tokenization and preprocessing, where we use SentencePiece.
  • Average Sentence Length: The average number of tokens per sentence in the dataset.
  • Word Overlap:
    $\dfrac{|T_{1}\cap T_{2}|}{|T_{1}|+|T_{2}|}$
    where $T_{1}$ and $T_{2}$ correspond to the unique tokens of two datasets.
  • Type-Token Ratio (TTR): The ratio of unique tokens to the total number of tokens in the dataset. This is given by:
    $\text{TTR}=\dfrac{|V|}{\sum_{i=1}^{N}|s_{i}|}$
    where $|V|$ is the number of unique tokens, and $\sum_{i=1}^{N}|s_{i}|$ is the total number of tokens in all $N$ sentences.
  • TTR Distance: The distance between the TTRs of two datasets, calculated as:
    $D_{\text{TTR}}=\left(1-\dfrac{\text{TTR}_{1}}{\text{TTR}_{2}}\right)^{2}$
    where $\text{TTR}_{1}$ and $\text{TTR}_{2}$ are the TTR values for two datasets.
  • Jensen-Shannon Divergence (JSD): Measures the divergence between two token distributions. For two token distributions $P$ and $Q$, the JSD is calculated as:
    $\text{JSD}(P,Q)=\dfrac{1}{2}\left[\text{KL}(P\parallel M)+\text{KL}(Q\parallel M)\right]$
    where $M=\dfrac{1}{2}(P+Q)$ and KL is the Kullback-Leibler divergence:
    $\text{KL}(P\parallel Q)=\sum_{x}P(x)\log\dfrac{P(x)}{Q(x)}$
  • TF-IDF Cosine Similarity: The cosine similarity between two datasets based on their TF-IDF representations. The cosine similarity is given by:
    Cosine $\text{Cosine Similarity}(A,B)=\dfrac{A\cdot B}{\|A\|\|B\|}$
    where $A$ and $B$ are the TF-IDF vectors of the two datasets.
  • Sentence-BERT Similarity: The cosine similarity between the mean sentence embeddings of two datasets obtained from Sentence-BERT Reimers and Gurevych (2019). It is computed as:
    Cosine $\text{Cosine Similarity}(E_{A},E_{B})=\dfrac{E_{A}\cdot E_{B}}{\|E_{A}\|\|E_{B}\|}$
    where $E_{A}$ and $E_{B}$ are the mean sentence embeddings for datasets $A$ and $B$, respectively.
Appendix B Experimental Details
@@
   FLoRes-200 Costa-jussà et al. (2022) afr, amh, arz, bak, bel, ceb, cym, deu, dik, ewe, Multi-domain
fao, fra, guj, hau, hne, hye, ibo, ind, jav, kan,
kat, kaz, khm, kin, kir, kor, lmo, ltz, mar, mri,
mya, oci, pan, plt, rus, sin, sna, snd, som, ssw,
tam, tat, tgl, tuk, wol, xho, vie, yor, zho, zul
@@
      Neulab TED Talk Qi et al. (2018) bel, deu, fra, hye, kat, kaz, kor, mar, mya, rus, TED
tam, vie, zho
      News Commentary deu, fra, rus, zho News
      News-test WMT22  Kocmi et al. (2022) deu, rus, zho News
      General-test WMT22  Kocmi et al. (2022) deu, fra, kaz, rus, zho News
      OPUS100 Zhang et al. (2020) afr, amh, bel, cym, deu, fra, guj, hau, hye, ibo, Multi-domain
ind, kan, kat, kaz, khm, kin, kir, kor, mar, mya,
oci, pan, rus, sin, tam, tat, tuk, vie, xho, yor,
zho, zul
      OPUS Bible afr, amh, ceb, deu, dik, ewe, fra, guj, hye, ind, Religion
      Christodouloupoulos and Steedman (2015) kan, kor, mar, mri, mya, plt, rus, sin, sna, som,
ssw, tam, tgl, vie, wol, xho, zho, zul
      OPUS OpenSubtitles afr, hye, kat, kor, sin, tgl, vie, zho Movies
@@
ind, jav, kan, kat, kaz, khm, kin, kir, kor, lmo,
ltz, mar, mri, mya, pan, sna, tat, tgl, tuk, vie,
yor, zul
      OPUS Tanzil Tiedemann (2012) amh, deu, fra, hau, ind, kor, rus, snd, som, tat, Religion
zho
      OPUS Gnome Tiedemann (2012) fao, mri, som, zho Technology
      OPUS GlobalVoices Tiedemann (2012) amh, kor, mya News
      OPUS Wikipedia Health Tiedemann (2012) fra, kor, rus, vie Health
      OPUS Antibiotic Tiedemann (2012) fra Health
      OPUS Tico Tiedemann (2012) fra, rus Health
      OPUS Vaccination Tiedemann (2012) fra Health
      OPUS Ubuntu Tiedemann (2012) vie, zho Technology
@@
      OPUS TLDR zho General
      UN rus Government
      lindat khresmoi Dušek et al. (2014) fra Health
@@
      WikiMatrix Schwenk et al. (2021) arz, bak, bel, ceb, fao, fra, ind, jav, kat, kaz, Wiki
kor, lmo, ltz, mar, oci, sin, tam, tat, tgl, vie,
zho
@@
Table 6: List of datasets under study covering 50 different languages. We only opt for 50 out of 500 languages available in the MT560 dataset, 50 out of 200 languages available in the FLoRes-200 dataset, and 8 out of 12 languages available in the NusaTranslation dataset.
B.1 Languages Under Study
We list all the languages used in the training from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets in Table 7 and Table 9, respectively. The language code follows ISO639-3 coding. All languages are also complemented by their rarity taxonomy based on Joshi et al. (2020) into two vitality classes: 0-2→low resource language (LRL), 3-4→mid resource language (MRL), and 5→high resource language (HRL). We also provide information about whether the language was part of the pre-trained M2M100 model dataset to highlight the model knowledge coverage.
B.2 Models
The details on the proxy models we use in MT experiments are as follows:
@@ B.2 Models
  • NLLB (No FT) Costa-jussà et al. (2022): [^4] NLLB (1.3B) is taken from https://github.com/facebookresearch/fairseq/tree/nllb. a pre-trained estimated model of NLLB-200 Distilled (1.3B) without any fine-tuning. We run the model in a zero-shot fashion.
The details on the proxy models we use in intent classification and slot filling experiments are as follows:
  • SmolLM (135M and 360M) Allal et al. (2024): [^5] SmolLM 135M and 360M are taken from https://huggingface.co/HuggingFaceTB/SmolLM-135M and https://huggingface.co/HuggingFaceTB/SmolLM-360M respectively a series of small LMs built on Cosmo-Corpus, a meticulously curated high-quality training dataset. SmolLM models have shown promising results when compared to other models in their size categories across various benchmarks testing common sense reasoning and world knowledge.
  • BLOOMZ (560M) Muennighoff et al. (2022): a decoder model fine-tuned from BLOOM Workshop et al. (2022) on a cross-lingual task mixture (xP3) and finds the resulting models capable of cross-lingual generalization to unseen tasks and languages.
Language Language Code Family Joshi Class Vitality Seen by M2M100 Covered by COMET-22
@@
Table 7: List of languages from the English-centric dataset on MT task, including their rarity category mapping, an indication of whether they are involved in the pre-training process for M2M100, and another indication of whether they are covered by COMET-22. Note that all languages are involved in the pre-training process for NLLB.
Language Language Code Family Joshi Class Vitality
Afrikaans afr indo-european 3 MRL
Amharic amh afro-asiatic 2 LRL
Arabic ara afro-asiatic 5 HRL
Azerbaijani aze turkic 1 LRL
Bengali ben indo-european 3 MRL
Welsh cym indo-european 1 LRL
Danish dan indo-european 3 MRL
German deu indo-european 5 HRL
Greek ell indo-european 3 MRL
English eng indo-european 5 HRL
Spanish spa indo-european 5 HRL
Persian fas indo-european 4 MRL
Finnish fin uralic 4 MRL
French fra indo-european 5 HRL
Hebrew heb afro-asiatic 3 MRL
Hindi hin indo-european 4 MRL
Hungarian hun uralic 4 MRL
Armenian hye indo-european 1 LRL
Indonesian ind austronesian 3 MRL
Icelandic isl indo-european 2 LRL
Italian ita indo-european 4 MRL
Japanese jpn japonic 5 HRL
Javanese jav austronesian 1 LRL
Georgian kat kartvelian 3 MRL
Khmer khm austro-asiatic 1 LRL
Kannada kan dravidian 1 LRL
Korean kor koreanic 4 MRL
Latvian lav indo-european 3 MRL
Malayalam mal dravidian 1 LRL
Mongolian mon mongolic 1 LRL
Malay msa austronesian 3 MRL
Burmese mya sino-tibetan 1 LRL
Norwegian Bokmål nob indo-european 1 LRL
Dutch nld indo-european 4 MRL
Polish pol indo-european 4 MRL
Portuguese por indo-european 4 MRL
Romanian ron indo-european 3 MRL
Russian rus indo-european 4 MRL
Slovenian slv indo-european 3 MRL
Albanian sqi indo-european 1 LRL
Swedish swe indo-european 4 MRL
Swahili swa niger-congo 2 LRL
Tamil tam dravidian 3 MRL
Telugu tel dravidian 1 LRL
Thai tha kra-dai 3 MRL
Tagalog tgl austronesian 3 MRL
Turkish tur turkic 4 MRL
Urdu urd indo-european 3 MRL
Vietnamese vie indo-european 4 MRL
Chinese (Simplified) zho sino-tibetan 5 HRL
Chinese (Traditional) zho sino-tibetan 5 HRL
Table 8: List of languages from the MASSIVE dataset for intent classification and slot filling, including their rarity category mapping.
B.3 Hyper-parameters
@@
Each fine-tuning and evaluation for LMs is done with an NVIDIA Tesla V100 32GB GPU.
The hyper-parameters used during fine-tuning from the English-centric and Many-to-Many Languages datasets for MT task are listed in Table 10, 11, 19, and 20 for SMaLL100, M2M100, NLLB, and Transformer models, respectively.
The common hyper-parameters used during fine-tuning for intent and slot classification tasks are listed in Table 12 for SmolLM-135M, SmolLM-360M, BLOOMZ-560M, Aya23-8B, and LLaMA3-8B respectively. The learning rate used for SmolLM-135M, SmolLM-360M, and BLOOMZ-560M is $1e^{-5}$, while the learning rate used for Aya23-8B and LLaMA3-8B is $1e^{-4}$. Since we are fine-tuning using SFT, we use the “default" template for SmolLM-135M, SmolLM-360M, and BLOOMZ-560M, while we use “cohere" for Aya23-8B and “llama3" for LLaMA3-8B. Additionally, we use LoRA for both Aya23-8B and LLaMA3-8B.
@@
Each regressor is trained on an AMD Ryzen Threadripper 2990WX with 128 GB of RAM and 16 threads. Regressors’ hyper-parameters used are provided in Table 21, 22, 23, 24, 25, and 26 for XGBoost, Poly2/Poly3, LGBM, and MF, respectively. These hyper-parameters were obtained based on the best cross-validation RMSE score on training using 10 folds.
@@
Table 9: List of languages from the Many-to-Many Languages dataset on MT task along with their rarity category mapping and an indication of whether they are included in the pre-training process for each respective model.
@@
Hyper-parameter Value
Cutoff Length 256
Preprocessing Num Workers 16
Train Batch Size 1
Gradient Accumulation Steps 2
Epochs 3
Learning Rate Scheduler cosine
Warmup Ratio 0.1
Eval Batch Size 1
Eval Steps 500
Table 12: List of common hyper-parameters used for fine-tuning SmolLM-135M, SmolLM-360M, BLOOMZ-560M, Aya23-8B, and LLaMA3-8B on intent and slot classification task.
B.4 Regressor Dataset Sizes
@@
Appendix C More Detailed Results
Models Random LOLO
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 7.69 $\pm$ 0.59 7.73 $\pm$ 0.08 9.20 $\pm$ 0.40 12.92 $\pm$ 0.54 9.16
Poly2 Khiu et al. (2024) 11.21 $\pm$ 0.49 16.23 $\pm$ 3.51 15.55 $\pm$ 0.00 43.02 $\pm$ 0.00 34.62
Poly3 Khiu et al. (2024) 11.00 $\pm$ 0.49 15.64 $\pm$ 2.59 62.29 $\pm$ 0.00 236.29 $\pm$ 0.00 66.97
LGBM Ye et al. (2021) 7.88 $\pm$ 0.65 8.15 $\pm$ 0.19 9.71 $\pm$ 0.17 12.81 $\pm$ 0.23 9.62
ProxyLM (Ours) with different proxy models
Transformer 4.68 $\pm$ 0.41 7.22 $\pm$ 0.25 6.18 $\pm$ 0.30 11.78 $\pm$ 0.50 7.31
SMaLL-100 4.07 $\pm$ 0.31 6.33 $\pm$ 0.11 4.59 $\pm$ 0.24 10.33 $\pm$ 0.41 6.04
SMaLL-100 (No FT) 5.27 $\pm$ 0.50 6.04 $\pm$ 0.32 6.28 $\pm$ 0.32 10.94 $\pm$ 0.50 6.95
Estimated Model (No FT) 5.23 $\pm$ 0.54 4.15 $\pm$ 0.24 6.18 $\pm$ 0.31 5.42 $\pm$ 0.27 5.38
Ensemble 3.21 $\pm$ 0.29 3.68 $\pm$ 0.36 3.74 $\pm$ 0.20 4.94 $\pm$ 0.29 4.01
Table 13: English-centric test results using spBLEU in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning" and the model inference is done in a zero-shot fashion. Avg represents the average of the results across the row. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models.
Models Random LOLO
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 0.0717 $\pm$ 0.0018 0.0556 $\pm$ 0.0038 0.1130 $\pm$ 0.0054 0.1007 $\pm$ 0.0046 0.0825
Poly2 Khiu et al. (2024) 0.1325 $\pm$ 0.0014 0.1503 $\pm$ 0.0019 0.1258 $\pm$ 0.0000 0.1176 $\pm$ 0.0000 0.1316
Poly3 Khiu et al. (2024) 0.1325 $\pm$ 0.0014 0.1520 $\pm$ 0.0019 0.1258 $\pm$ 0.0000 0.1176 $\pm$ 0.0000 0.1320
LGBM Ye et al. (2021) 0.0752 $\pm$ 0.0025 $0$$\pm$ 0.0033 0.1147 $\pm$ 0.0000 0.1019 $\pm$ 0.0000 0.0880
ProxyLM (Ours) with different proxy models
Transformer 0.0543 $\pm$ 0.0009 0.0492 $\pm$ 0.0041 0.0837 $\pm$ 0.0042 0.0957 $\pm$ 0.0031 0.0707
SMaLL-100 0.0328 $\pm$ 0.0012 0.0431 $\pm$ 0.0033 0.0547 $\pm$ 0.0027 0.0824 $\pm$ 0.0030 0.0533
SMaLL-100 (No FT) 0.0454 $\pm$ 0.0024 0.0393 $\pm$ 0.0033 0.0778 $\pm$ 0.0030 0.0967 $\pm$ 0.0034 0.0648
Estimated Model (No FT) 0.0460 $\pm$ 0.0017 0.0348 $\pm$ 0.0016 0.0777 $\pm$ 0.0034 0.0607 $\pm$ 0.0034 0.0548
Ensemble 0.0289 $\pm$ 0.0012 0.0293 $\pm$ 0.0013 0.0454 $\pm$ 0.0019 0.0613 $\pm$ 0.0031 0.0412
Table 14: English-centric MT test results using COMET-22 in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning" and the model inference is done in a zero-shot fashion. There are no results for the Unseen setting since COMET-22 score does not cover the unseen languages to M2M100. Furthermore, all languages are covered by NLLB in English-centric dataset. Avg represents the average of the results across the row. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models, the detailed breakdown of this result can be seen in Section C in the Appendix.
Models Random LOLO
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 2.45 $\pm$ 0.30 1.11 $\pm$ 0.06 7.83 $\pm$ 0.23 8.28 $\pm$ 0.31 4.94
LGBM  Ye et al. (2021) 2.66 $\pm$ 0.22 1.76 $\pm$ 0.29 7.91 $\pm$ 0.01 8.06 $\pm$ 0.00 5.10
ProxyLM (Ours) with different proxy models
Transformer 2.56 $\pm$ 0.43 1.70 $\pm$ 0.20 5.65 $\pm$ 0.23 6.24 $\pm$ 0.34 4.04
SMaLL-100 2.56 $\pm$ 0.33 1.65 $\pm$ 0.44 4.85 $\pm$ 0.36 5.14 $\pm$ 0.46 3.55
Estimated Model (No FT) 2.38 $\pm$ 0.36 1.27 $\pm$ 0.03 5.10 $\pm$ 0.28 5.50 $\pm$ 0.26 3.56
Ensemble 2.41 $\pm$ 0.28 1.56 $\pm$ 0.35 3.73 $\pm$ 0.23 3.79 $\pm$ 0.19 2.90
Table 15: Many-to-Many Languages test results in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning". Avg represents the average of the results across the row. The reported results are experiments using XGBoost regressor. Ensemble denotes combining all four proxy models.
Models Unseen Cross-Dataset
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 8.26 $\pm$ 0.53 12.90 $\pm$ 1.01 12.02 $\pm$ 1.02
Poly2  Khiu et al. (2024) 9.51 $\pm$ 0.00 11.02 $\pm$ 0.00 8.97 $\pm$ 0.00
Poly3  Khiu et al. (2024) 9.64 $\pm$ 0.00 11.06 $\pm$ 0.00 10.98 $\pm$ 0.00
LGBM  Ye et al. (2021) 9.56 $\pm$ 0.97 9.34 $\pm$ 0.00 10.58 $\pm$ 0.00
ProxyLM (Ours) with different proxy models
Transformer 6.71 $\pm$ 0.37 9.31 $\pm$ 0.00 9.70 $\pm$ 0.00
SMaLL-100 4.87 $\pm$ 0.34 4.52 $\pm$ 0.05 6.87 $\pm$ 0.11
SMaLL-100 (No FT) 6.22 $\pm$ 0.38 8.26 $\pm$ 0.59 9.50 $\pm$ 0.04
Estimated Model (No FT) 5.90 $\pm$ 0.40 9.95 $\pm$ 0.40 7.20 $\pm$ 0.24
Ensemble 4.48 $\pm$ 0.23 4.35 $\pm$ 0.02 5.03 $\pm$ 0.48
Table 16: Unseen and Cross-Dataset MT test results on English-centric dataset in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning". We only show M2M100 results for the Unseen setting since NLLB covers all languages in the English-centric dataset. The reported results for the Unseen setting use XGBoost, while the Cross-Dataset experiments use LGBM. Ensemble denotes combining all four proxy models.
Models Random LOLO
Aya-23$\downarrow$ LLaMA3$\downarrow$ Aya-23$\downarrow$ LLaMA3$\downarrow$ Avg.
NLPerf Xia et al. (2020) with different regressors
XGBoost 0.0761 $\pm$ 0.0008 0.0191 $\pm$ 0.0008 0.1573 $\pm$ 0.0039 0.0581 $\pm$ 0.0021 0.0777
Poly2 Khiu et al. (2024) 0.1996 $\pm$ 0.0017 0.0979 $\pm$ 0.0033 0.2075 $\pm$ 0.0000 0.0918 $\pm$ 0.0000 0.1492
Poly3 Khiu et al. (2024) 0.1990 $\pm$ 0.0011 0.0969 $\pm$ 0.0015 0.2191 $\pm$ 0.0000 0.0925 $\pm$ 0.0000 0.1519
LGBM Ye et al. (2021) 0.0839 $\pm$ 0.0030 0.0198 $\pm$ 0.0023 0.1545 $\pm$ 0.0000 0.0558 $\pm$ 0.0006 0.0785
ProxyLM (Ours) with different proxy models
SmolLM (135M) 0.0676 $\pm$ 0.0013 0.0171 $\pm$ 0.0008 0.1273 $\pm$ 0.0033 0.0455 $\pm$ 0.0020 0.0644
SmolLM (360M) 0.0604 $\pm$ 0.0015 0.0157 $\pm$ 0.0003 0.1118 $\pm$ 0.0031 0.0441 $\pm$ 0.0017 0.0580
BLOOMZ (560M) 0.0692 $\pm$ 0.0025 0.0179 $\pm$ 0.0009 0.1283 $\pm$ 0.0038 0.0482 $\pm$ 0.0023 0.0659
Ensemble 0.0609 $\pm$ 0.0010 0.0164 $\pm$ 0.0008 0.1112 $\pm$ 0.0032 0.0442 $\pm$ 0.0019 0.0582
Table 17: Intent classification results using accuracy in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. Avg represents the average of the results across the row. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models.
Models Random LOLO
Aya-23$\downarrow$ LLaMA3$\downarrow$ Aya-23$\downarrow$ LLaMA3$\downarrow$ Avg.
NLPerf Xia et al. (2020) with different regressors
XGBoost 0.0693 $\pm$ 0.0009 0.0548 $\pm$ 0.0034 0.1219 $\pm$ 0.0035 0.1093 $\pm$ 0.0027 0.0888
Poly2 Khiu et al. (2024) 0.1396 $\pm$ 0.0008 0.1412 $\pm$ 0.0017 0.1418 $\pm$ 0.0000 0.1414 $\pm$ 0.0000 0.1410
Poly3 Khiu et al. (2024) 0.1393 $\pm$ 0.0008 0.1401 $\pm$ 0.0017 0.1448 $\pm$ 0.0000 0.1413 $\pm$ 0.0000 0.1414
LGBM Ye et al. (2021) 0.0692 $\pm$ 0.0017 0.0557 $\pm$ 0.0029 0.1218 $\pm$ 0.0000 0.1152 $\pm$ 0.0000 0.0905
ProxyLM (Ours) with different proxy models
SmolLM (135M) 0.0618 $\pm$ 0.0018 0.0538 $\pm$ 0.0022 0.1004 $\pm$ 0.0028 0.0953 $\pm$ 0.0024 0.0778
SmolLM (360M) 0.0562 $\pm$ 0.0011 0.0506 $\pm$ 0.0023 0.0844 $\pm$ 0.0024 0.0868 $\pm$ 0.0018 0.0695
BLOOMZ (560M) 0.0618 $\pm$ 0.0027 0.0540 $\pm$ 0.0019 0.1023 $\pm$ 0.0033 0.0995 $\pm$ 0.0028 0.0794
Ensemble 0.0561 $\pm$ 0.0016 0.0508 $\pm$ 0.0021 0.0830 $\pm$ 0.0022 0.0884 $\pm$ 0.0025 0.0696
Table 18: Slot filling results using micro-F1 score in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. Avg represents the average of the results across the row. The reported results use LGBM as the regressor. Ensemble denotes combining all four proxy models.
We provide detailed results for Table 1, 2, and 2 by providing the standard errors of the predictions. The mapping of vitality, Joshi class, and language family follows the classifications in Table 7 and 9. The mapping of all languages in Table 5 until 7.
Appendix D Feature Importances